OpenStudy (anonymous):

solve: 2y^(1/2) + 0.2y not sure how to.. trying 2.2*y^(3/2) is that right??

OpenStudy (anonymous):

\[2\sqrt{y}+0.2y=0\]???

OpenStudy (anonymous):

I need to solve integral: 2pi*y*[(2rooty-(-0.2y)]dy

OpenStudy (anonymous):

oh sorry :)

OpenStudy (anonymous):

mhm

hartnn (hartnn):

u have to solve \(\large \int2\pi y(2\sqrt y+0.2y)dy=\int 4piy^{3/2}dy+\int0.4\pi y^2dy \) did u get this step ? can u solve further ?

hartnn (hartnn):

just take constants out of integral and use \(\huge \int x^ndx=\frac{x^{n+1}}{n+1}+c\)

OpenStudy (anonymous):

ive been doing that, getting it wrong

OpenStudy (anonymous):

err i did a diff. problem and got it..was doing the #s on calculator wrong..ty

hartnn (hartnn):

so u got the correct answer finally?

OpenStudy (anonymous):

damn it im not getting it..

hartnn (hartnn):

(8/5)pi y^(5/2) +(0.4/3)pi y^3 u getting this ?

OpenStudy (anonymous):

damn it i was off by decimal points..b/c its the software on computer lol

OpenStudy (anonymous):

i got it correct now..

hartnn (hartnn):

great :)